OpenStudy (anonymous):

dividing radicals helpppppp square root of 5 divided by 5 + square root of 2 2 - square root of 3divided by

OpenStudy (anonymous):

\[\sqrt{5}\div5+\sqrt{2}\]

OpenStudy (anonymous):

\[2-\sqrt{3}\div-2-\sqrt{5}\]

OpenStudy (anonymous):

that would give you 1.861427128....

OpenStudy (anonymous):

but how did u get that

OpenStudy (anonymous):

the first one..gave you that

OpenStudy (anonymous):

I have a texas ti-84 plues

OpenStudy (anonymous):

i mean step by step

OpenStudy (anonymous):

yea so do i

OpenStudy (anonymous):

put it in you cal.

OpenStudy (anonymous):

to put in the first problem

OpenStudy (anonymous):

i know but i need to show the work

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

so in you cal. do square root of 5

OpenStudy (anonymous):

i need to show the work on papaer

OpenStudy (anonymous):

I dont know how to show it on paper...

OpenStudy (anonymous):

ur teacher must be trippin

OpenStudy (anonymous):

thnkss anyway

OpenStudy (anonymous):

i guess

OpenStudy (radar):

\[\sqrt{5}\over5+\sqrt{2}\]

OpenStudy (radar):

Step 1 multiply both the numerator by the conjugate of the denominator. In this problem the conjugate is:\[5-\sqrt{2}\] so your problem would now look like this:\[\sqrt{5}(5-\sqrt{2})\over(5+\sqrt{2}) (5-\sqrt{2}\] do you follow so far?

OpenStudy (anonymous):

yess i do

OpenStudy (anonymous):

r u there

OpenStudy (radar):

Our fraction now becomes:\[5\sqrt{5}-\sqrt{10}\over25-2\] or\[\sqrt{5}(5-\sqrt{2})\over23\]\[2.236(5-1.414)\over23\]\[(2.236)(3.585)\over23\]\[8.01606\over23\] .3485

OpenStudy (anonymous):

thnk u theres another problem

OpenStudy (radar):

You could, rather than rationalize the denominator just use your calculator. Lets do that on the second one.

OpenStudy (anonymous):

ok thnks

OpenStudy (radar):

\[2-\sqrt{3}\over-2-\sqrt{5}\] \[2-1.732\over-2-2.236\]\[.268\over-4.236\] -.06326

OpenStudy (radar):

There is an error somewhere, the answer is negative, but larger negative than that!